Write the following equations in slope-intercept form:

a) y + 2 = -6(x – 1)
b) 1 – 2x – y = 0


Answer by tutorcecilia(2152)

a) y + 2 = -6(x – 1) [Simplify]
y + 2 = -6x + 6 [Solve for y]
y + 2 -2 = -6x + 6 - 2
y = -6x + 4

.

b) 1 – 2x – y = 0 [Solve for y]
1-2x=y [Switch the order]
-2x-1=y or y=-2x-1